# RSAdapter: Adapting Multimodal Models for Remote Sensing Visual Question Answering

### v1

@@ Abstract
In recent years, with the rapid advancement of transformer models, transformer-based multimodal architectures have found wide application in various downstream tasks, including but not limited to Image Captioning, Visual Question Answering (VQA), and Image-Text Generation. However, contemporary approaches to Remote Sensing (RS) VQA often involve resource-intensive techniques, such as full fine-tuning of large models or the extraction of image-text features from pre-trained multimodal models, followed by modality fusion using decoders. These approaches demand significant computational resources and time, and a considerable number of trainable parameters are introduced. To address these challenges, we introduce a novel method known as RSAdapter, which prioritizes runtime and parameter efficiency. RSAdapter comprises two key components: the Parallel Adapter and an additional linear transformation layer inserted after each fully connected (FC) layer within the Adapter. This approach not only improves adaptation to pre-trained multimodal models but also allows the parameters of the linear transformation layer to be integrated into the preceding FC layers during inference, reducing inference costs. To demonstrate the effectiveness of RSAdapter, we conduct an extensive series of experiments using three distinct RS-VQA datasets and achieve state-of-the-art results on all three datasets. The code for RSAdapter will be available online at https://github.com/Y-D-Wang/RSAdapter.
@@ I Introduction
VQA has recently gained significant attention in the field of Earth Science and Remote Sensing (RS), particularly with the introduction of several new datasets [14, 15, 16]. This makes it a current and cutting-edge research topic in these domains. Lobery et al. [14] were among the first to introduce a simple method where features were extracted separately from images using CNN and from text using LSTM. These features were then combined through element-wise multiplication to obtain a joint representation. Subsequent works [17, 18] in this field explored the use of visual features at different levels to enhance the understanding of image content, thereby improving the model’s ability to learn from both images and text. The methods mentioned above are all based on traditional CNN and RNN approaches. Similarly, models based on transformers have also been proposed in this context. Bazi et al. [19] utilized the CLIP [20] model to extract image and text features. These features were then separately fed into two parallel decoder transformers to learn a joint representation. Siebert et al. [21] fine-tuned datasets using the VisualBERT [7] model.
@@
  • In the domain of natural images, it has been demonstrated that well pre-trained multimodal models can offer strong transferability to downstream VQA tasks through fine-tuning. However, objects in remote sensing images are often more challenging to distinguish than those in natural images. Currently, transformer-based methods have not yielded the best results in RS-VQA. Therefore, there are doubts about whether pretrained transformer models still exhibit strong transferability on RS-VQA tasks.
  • The majority of methods require updating all parameters of the pre-trained model to adapt to the RS-VQA tasks and achieve satisfactory performance. Due to the increasing size of pre-trained models, the computational cost has also grown significantly.
  • There is still debate over whether it is necessary to perform full fine-tuning on pre-trained multimodal models, given their demonstrated strong transferability to downstream tasks. For data-limited RS-VQA tasks, pre-trained models with a large number of parameters may lose their generalization ability after fine-tuning.
To overcome the limitations mentioned above, we propose a novel approach, namely RSAdapter, to efficiently fine-tune pre-trained multimodal models, enabling them to better adapt to RS-VQA tasks. Specifically, we start with the ViLT model [22] as our backbone, and then we compare the performance differences between inserting RSAdapter in parallel next to the MSA and MLP components separately. The results consistently show that RSAdapter inserted next to the MLP component performed better. Finally, we simultaneously insert RSAdapter next to MSA and MLP components while adding corresponding scaling layers to control the contribution of each RSAdapter to the model effectively. To validate the effectiveness of RSAdapter, we conduct comprehensive experiments on three different RS-VQA datasets, and the results indicated that when freezing all parameters of the pre-trained multimodal model and only updating the newly added lightweight RSAdapter, our proposed method achieves better results than previous state-of-the-art methods. The main contributions of this paper are summarized as follows:
@@
    In contrast to conventional methods for RS-VQA, our approach achieves impressive performance without the reliance on region-based features or complex feature extractors for visual and textual embeddings.
@@
    We introduce a novel re-parameterization method, RSAdapter, which can achieve better performance without increasing the number of parameters in the inference phase. Compared to RSVQA, our inference time can be reduced by 43% on the RS-LR dataset.
  4. 4.
@@ II-A Vision-Language Transformer Models
Currently, most vision-language models use a single-stream approach, which includes only one transformer stack together. They concatenate visual tokens and textual tokens as the input to the transformer. Due to its simplicity, it significantly simplifies the architecture compared to the dual-streams transformers, achieving comparable performance with fewer parameters. Additionally, it offers ease of scalability, as we can handle other modalities’ questions, such as video and audio, by simply concatenating tokens from other modalities. ViLT [22], PixelBERT [26], VisualBERT [7] all belong to the category of single-stream transformers. They all use BERT [4] to extract text embeddings, and both PixelBERT and VisualBERT employ CNNs as feature extractors for the image modality. After specific processing, the image features are concatenated with text tokens and fed into the transformer. What sets ViLT apart from other models is its use of a straightforward linear projection to obtain image embeddings. This simplifies the structure of ViLT, leading to increased runtime and parameter efficiency. It is precisely this characteristic that led us to choose ViLT as our backbone model in this paper.
@@ II-C Remote Sensing VQA
Lobery et al. [14] were the first to introduce VQA to the RS field. Their work initially released two new datasets, referred to as RS-VQA Low Resolution and RS-VQA High Resolution. Subsequently, they proposed a simple joint approach, which involved extracting features from both images and text data using CNN and LSTM separately, followed by computing the dot product of image-text features and feeding them into an MLP for answer prediction. Later, they released a large-scale VQA dataset [16] dedicated to remote sensing, containing close to 15 million samples.
@@ III-A Preliminary
where $P_{t}\in\mathbb{R}^{(n_{t}+1)\times d}$ and $P_{v}\in\mathbb{R}^{(n_{v}+1)\times d}$ are the positional embeddings for text embeddings and image embeddings. Both are combined with modality type embeddings, where 00 and $1$ represent text $t_{type}$ and image $v_{type}$, respectively. They are then concatenated to create a unified token denoted as $X_{0}$, which can be formulated by,
@@
$\text{Head${}_{i}$}(X)=\text{Softmax}(\frac{(XW_{i}^{Q})(XW_{i}^{K^{\mathrm{T}}})}{\sqrt{d_{k}}}(XW_{i}^{V}))$ (7)
where $\text{Head${}_{i}$}(X)$, $W_{i}^{Q}\in\mathbb{R}^{d\times\frac{d}{h}}$, $W_{i}^{K}\in\mathbb{R}^{d\times\frac{d}{h}}$ and $W_{i}^{V}\in\mathbb{R}^{d\times\frac{d}{h}}$ are the $i$-th head dot-product attention and weight matrices. $W^{O}\in\mathbb{R}^{d\times d}$ is the weight matrix for the output transformation. MLP can be formulated by
@@ III-C RSAdapter
Here, we introduce a novel re-parameterization method that allows us to achieve better performance during inference without incurring additional costs, building upon the foundation of the parallel Adapter. We first add a linear transformation after each Linear layer in the parallel Adapter. In this case, the Eq. (12) can be rewritten as
@@ IV-B Experimental Settings and Implementation Details
[…]
@@ IV-D1 Effectiveness of Components
In Table V, we present a detailed comparison of different components of RSAdapter and other PEFT methods, including Bitfit [34], Adapter [27], and Lora [30]. Additionally, we report the results of linear probing and full fine-tuning for reference. To analyze the impact of different RSAdapter insertion positions, we provide two variants: RSAdapter(MSA), inserted in parallel alongside MSA, and RSAdapter(MLP), inserted in parallel alongside MLP. From the table, it is evident that the full RSAdapter consistently achieves the best results across all three different test sets. Notably, RSAdapter(MLP) exhibits varying degrees of improvement over RSAdapter(MSA) across all datasets. This suggests that during the model’s efficient fine-tuning stage, tuning the MLP tends to be more beneficial than tuning the MSA. When compared to other PEFT methods, the full RSAdapter significantly outperforms them when tuning an equivalent number of parameters. It is worth noting that, in the majority of datasets, the results of the Adapter method are superior to the other two methods, further confirming the effectiveness of the Adapter approach. Compared to full fine-tuning of the entire model, our approach achieves comparable results with only a small fraction of parameters tuned and, in some datasets, even exhibits slight improvements.
IV-D2 Different Backbone Models
In further validation of the applicability of our method, we conduct additional experiments using two unimodal large models, ViT [41] and BERT [4]. Since ViT and BERT can only handle one modality of input each, we perform element-wise multiplication between the image features and text features obtained from these two models. Subsequently, we feed this combined image-text feature into the classification layer to obtain our answers. From Table VI, it is evident that our method, with only half the parameter updates compared to ViT+BERT, outperforms it on three different datasets when using the ViLT model. This suggests that, even with a smaller parameter count in the multimodal backbone model, our method performs better than using unimodal backbone models.
TABLE VI: Performance comparision using different Backbone on three different data sets. All values are reported as percentage (%).
@@ IV-D2 Different Backbone Models
IV-D3 Data Efficiency
TABLE VII: Data efficiency results on three different data sets. Results are reported as percentage (%). Best results are highlighted in bold.
@@ IV-D3 Data Efficiency
To assess the data efficiency of our method, we analyze RSAdapter’s performance under conditions of limited training data. We employ dataset subsets representing proportions of 0.1, 0.2, and 0.5 and report the results of all three datasets in Table VII. Remarkably, even with very limited data, we are able to achieve acceptable results. Notably, our method can outperform other models trained on the complete dataset, even when using only half of the data (e.g., achieving a 2.19% increase for AA and a 1.21% increase for OA on the RSIVQA dataset). In Fig. 3, we compare the results of RSAdapter and Bi-Modal when using only 10% and 20% of the data on LR and HR datasets. RSAdapter outperforms Bi-Modal on both LR and HR datasets, except for the AA on 10% of LR. This discrepancy is due to the class imbalance in the LR dataset, where the Rural/Urban category has a significant impact on the final AA. In Fig. 4, we also present the results of RSAdapter on the RSIVQA dataset across various training subsets, different categories, AA and OA. In conclusion, our method demonstrates the ability to achieve satisfactory results even with limited data.
@@
IV-D4 Different Bottleneck Dimension Size
TABLE VIII: Effect of bottleneck size $d^{\prime}$ of RSAdapters on the LR test set. Results are reported as percentage (%). Best results are highlighted in bold.
@@ IV-D4 Different Bottleneck Dimension Size
The bottleneck dimension size of RSAdapters is a crucial hyperparameter that affects both the number of model parameters and tunable parameters. To investigate its impact on model performance, we select $d^{\prime}$ from the set ${32,64,128,192,256,384}$. Table VIII presents the results of the model on the LR test set for different values of $d^{\prime}$. Upon analyzing the results, we observe that the model performs best when $d^{\prime}$ is set to 192. However, as $d^{\prime}$ increases, the model’s parameter count also rises, but there is no corresponding improvement in performance. This suggests that the increase in $d^{\prime}$ may lead to increased optimization complexity, resulting in worse performance. Therefore, selecting the appropriate $d^{\prime}$ is crucial.
IV-D5 Different Position of RSAdapters
In addition to the experiments mentioned above, we also conduct experiments to assess the impact of the insertion positions of RSAdapter, using the same tunable parameters for a fair comparison. Table IX provides results for different insertion positions: ’Top’ refers to inserting RSAdapter into the last 6 layers of the transformer, ’Bottom’ refers to inserting RSAdapter into the first 6 layers of the transformer, while ’Even’ and ’Odd’ stand for inserting RSAdapter into the even and odd layers of the transformer, respectively. Remarkably, when RSAdapter is inserted into the lower layers, the results consistently outperform those when inserted into the upper layers. This could be attributed to the fact that there may be slight differences in low-level features between natural images and remote sensing images, leading to significant variations in results when fine-tuning a model pre-trained on natural images for remote sensing images.
TABLE IX: Effect of different positions of RSAdapters on the LR test set. Results are reported as percentage (%). Best results are highlighted in bold.

### v2

@@ Abstract
In recent years, with the rapid advancement of transformer models, transformer-based multimodal architectures have found wide application in various downstream tasks, including but not limited to Image Captioning, Visual Question Answering (VQA), and Image-Text Generation. However, contemporary approaches to Remote Sensing (RS) VQA often involve resource-intensive techniques, such as full fine-tuning of large models or the extraction of image-text features from pre-trained multimodal models, followed by modality fusion using decoders. These approaches demand significant computational resources and time, and a considerable number of trainable parameters are introduced. To address these challenges, we introduce a novel method known as RSAdapter, which prioritizes runtime and parameter efficiency. RSAdapter comprises two key components: the Parallel Adapter and an additional linear transformation layer inserted after each fully connected (FC) layer within the Adapter. This approach not only improves adaptation to pre-trained multimodal models but also allows the parameters of the linear transformation layer to be integrated into the preceding FC layers during inference, reducing inference costs. To demonstrate the effectiveness of RSAdapter, we conduct an extensive series of experiments using three distinct RS-VQA datasets and achieve state-of-the-art results on all three datasets. The code for RSAdapter is available online at https://github.com/Y-D-Wang/RSAdapter.
@@ I Introduction
VQA has recently gained significant attention in the field of Earth Science and Remote Sensing (RS), particularly with the introduction of several new datasets [14, 15, 16]. This makes it a current and cutting-edge research topic in these domains. Lobry et al. [14] were among the first to introduce a simple method where features were extracted separately from images using CNN and from text using LSTM. These features were then combined through element-wise multiplication to obtain a joint representation. Subsequent works [17, 18] in this field explored the use of visual features at different levels to enhance the understanding of image content, thereby improving the model’s ability to learn from both images and text. The methods mentioned above are all based on traditional CNN and RNN approaches. Similarly, models based on transformers have also been proposed in this context. Bazi et al. [19] utilized the CLIP [20] model to extract image and text features. These features were then separately fed into two parallel decoder transformers to learn a joint representation. Siebert et al. [21] jointly fed processed image and text features into VisualBert [7] for better understanding of image-text features. Hackel et al. [22] presented a lightweight transformer-based VQA model in remote sensing and also published a PyTorch-based library for rapid development of Image-Language Models.
@@
  • In the domain of natural images, it has been demonstrated that well pre-trained multimodal models can offer strong transferability to downstream VQA tasks through fine-tuning. RS images typically exhibit complex backgrounds, high similarity between different objects, and large object size variations, making them more challenging to process than natural images. Therefore, there are doubts on whether pretrained transformer models used for natural images still exhibit strong transferability on RS-VQA tasks.
  • The majority of methods require updating all parameters of the pre-trained model to adapt to the RS-VQA tasks and achieve satisfactory performance. As the scale of pre-trained large models rapidly expands, for example, the Llama2 model [23] with 70B parameters, saving computational costs has remained a significant challenge.
  • There is still debate over whether it is necessary to perform full fine-tuning on pre-trained multimodal models, given their demonstrated strong transferability to downstream tasks. Directly applying full fine-tuning techniques of pre-tranined large models on data-limited RS-VQA tasks may lead to overfitting, which results in inferior performance, instability, and reduced generalization ability. This, in turn, affects the effectiveness of utilizing pre-trained models.
To overcome the limitations mentioned above, we propose a novel approach, namely RSAdapter, to efficiently fine-tune pre-trained multimodal models, enabling them to better adapt to RS-VQA tasks. Specifically, we start with the ViLT model [24] as our backbone, and then we compare the performance differences between inserting RSAdapter in parallel next to the multiheaded self attention (MSA) and feedforward network (MLP) components separately. The results consistently show that RSAdapter inserted next to the MLP component performed better. Finally, we simultaneously insert RSAdapter next to MSA and MLP components while adding corresponding scaling layers to control the contribution of each RSAdapter to the model effectively. To validate the effectiveness of RSAdapter, we conduct comprehensive experiments on three different RS-VQA datasets, and the results indicated that when freezing all parameters of the pre-trained multimodal model and only updating the newly added lightweight RSAdapter, our proposed method achieves better results than previous state-of-the-art methods. The main contributions of this paper are summarized as follows:
@@
    In contrast to conventional methods for RS-VQA, our approach achieves impressive performance without the reliance on region-based features or complex feature extractors for visual and textual embeddings. Moreover, we introduce a novel re-parameterization method, RSAdapter, which can achieve better performance without increasing the number of parameters in the inference phase. Compared to RSVQA, our inference time can be reduced by 43% on the RS-LR dataset.
@@ II-A Vision-Language Transformer Models
Currently, most vision-language models use a single-stream approach, which includes only one transformer stack together. They concatenate visual tokens and textual tokens as the input to the transformer. It significantly simplifies the architecture compared to the dual-streams transformers, achieving comparable performance with fewer parameters. Additionally, it offers ease of scalability, as we can handle other modalities’ questions, such as video and audio, by simply concatenating tokens from other modalities. ViLT [24], PixelBERT [28], VisualBERT [7] all belong to the category of single-stream transformers. They all use BERT [4] to extract text embeddings, and both PixelBERT and VisualBERT employ CNNs as feature extractors for the image modality. After specific processing, the image features are concatenated with text tokens and fed into the transformer. What sets ViLT apart from other models is its use of a straightforward linear projection to obtain image embeddings. This simplifies the structure of ViLT, leading to increased runtime and parameter efficiency. It is precisely this characteristic that led us to choose ViLT as our backbone model in this paper.
@@ II-C Remote Sensing VQA
Lobry et al. [14] were the first to introduce VQA to the RS field. Their work initially released two new datasets, referred to as RS-VQA Low Resolution and RS-VQA High Resolution. Subsequently, they proposed a simple joint approach, which involved extracting features from both images and text data using CNN and LSTM separately, followed by computing the dot product of image-text features and feeding them into an MLP for answer prediction. Later, they released a large-scale VQA dataset [16] dedicated to remote sensing, containing close to 15 million samples.
@@ III-A Preliminary
where $P_{t}\in\mathbb{R}^{(n_{t}+1)\times d}$ and $P_{v}\in\mathbb{R}^{(n_{v}+1)\times d}$ are the positional embeddings for text embeddings and image embeddings. To enhance the model’s ability to discern the relationship between images and text, corresponding modal-type embeddings are added to text and image embeddings. Text modal-type embedding $t_{type}$ is assigned 00, and image modal-type embedding $v_{type}$ is assigned $1$, enabling the model to differentiate between images and text. They are then concatenated to create a unified token denoted as $X_{0}$, which can be formulated by,
@@
[…]
@@ III-C RSAdapter
Re-parameterization techniques have been widely applied in the field of CV [46, 47]. Here, we introduce a novel re-parameterization method that allows us to achieve better performance during inference without incurring additional costs, building upon the foundation of the parallel Adapter. We first add a linear transformation after each Linear layer in the parallel Adapter. In this case, the Eq. (12) can be rewritten as
@@ IV-B Experimental Settings and Implementation Details
In the training phase, we set the maximum number of iterations to 50 epochs for the LR and RSIVQA datasets and 20 epochs for the HR dataset. We use a batch size of 64. For the LR and RSIVQA datasets, we employ a learning rate of 1e-3 for the first 4 epochs to warm up our model, after which we optimize the model using a learning rate of 1e-5. In the case of the HR dataset, the warm-up learning rate is set to 1e-4, and the normal learning rate is 1e-6. We utilize the Adam optimizer [49] for model optimization.
@@ IV-C5 SHRNet
In addition, we have separately reported the results of the question-only model in Table I to Table IV. This model masks all input images as 0 to predict answers, demonstrating the performance of the model when only textual information is provided. The final results show that the question-only model exhibits lower accuracy compared to the image-text model.
@@ IV-D1 Effectiveness of Components
In Table V, we present a detailed comparison of different components of RSAdapter and other PEFT methods, including Bitfit [36], Adapter [29], and Lora [32]. Additionally, we report the results of linear probing and full fine-tuning for reference. The setup for linear probing is as follows: all parameters of the ViLT model are frozen, and only the parameters of the final classifier are updated.
To analyze the impact of different RSAdapter insertion positions, we provide two variants: RSAdapter(MSA), inserted in parallel alongside MSA, and RSAdapter(MLP), inserted in parallel alongside MLP. From the table, it is evident that the full RSAdapter consistently achieves the best results across all three different test sets. Notably, RSAdapter(MLP) exhibits varying degrees of improvement over RSAdapter(MSA) across all datasets. This suggests that during the model’s efficient fine-tuning stage, tuning the MLP tends to be more beneficial than tuning the MSA. When compared to other PEFT methods, the full RSAdapter significantly outperforms them when tuning an equivalent number of parameters. It is worth noting that, in the majority of datasets, the results of the Adapter method are superior to the other two methods, further confirming the effectiveness of the Adapter approach. Compared to full fine-tuning of the entire model, our approach achieves comparable results with only a small fraction of parameters tuned and, in some datasets, even exhibits slight improvements.
IV-D2 Impact of skip connection in RSAdapter
Table VI presents a comparison of the results of RSAdapter with skip connection removed and RSAdapter with skip connection retained in different scenarios. We observe that removing the skip connection consistently results in better AA and OA in all scenarios. It is worth noting that compared to RSAdapter with skip connection added only alongside MLP, adding skip connection alongside both MSA and MLP in RSAdapter simultaneously can lead to a decrease in performance. Therefore, we have decided to remove the skip connection in RSAdapter.
TABLE VI: Performance comparision between RSAdapter with and without skip connection on the LR test set. All values are reported as percentage (%).
Type Param (M) Tunable Param (M) Average Accuracy Overall Accuracy
MSA w/ sc 116 4.8 86.95 85.94
MSA w/o sc 116 4.8 86.75 86.33
MLP w/ sc 116 4.8 87.07 86.47
MLP w/o sc 116 4.8 87.63 86.73
MSA+MLP w/ sc 120 8.4 86.58 86.38
MSA+MLP w/o sc 120 8.4 87.78 87.14
Note: sc indicates skip connection.
IV-D3 Different Backbone Models
In further validation of the applicability of our method, we conduct additional experiments using two unimodal large models, ViT [43] and BERT [4]. Since ViT and BERT can only handle one modality of input each, we perform element-wise multiplication between the image features and text features obtained from these two models. Subsequently, we feed this combined image-text feature into the classification layer to obtain our answers. From Table VII, it is evident that our method, with only half the parameter updates compared to ViT+BERT, outperforms it on three different datasets when using the ViLT model. This suggests that, even with a smaller parameter count in the multimodal backbone model, our method performs better than using unimodal backbone models.
TABLE VII: Performance comparision using different Backbone on three different data sets. All values are reported as percentage (%).
@@ IV-D3 Different Backbone Models
IV-D4 Data Efficiency
TABLE VIII: Data efficiency results on three different data sets. Results are reported as percentage (%). Best results are highlighted in bold.
@@ IV-D4 Data Efficiency
To assess the data efficiency of our method, we analyze RSAdapter’s performance under conditions of limited training data. We employ dataset subsets representing proportions of 0.1, 0.2, and 0.5 and report the results of all three datasets in Table VIII. Remarkably, even with very limited data, we are able to achieve acceptable results. Notably, our method can outperform other models trained on the complete dataset, even when using only half of the data (e.g., achieving a 2.19% increase for AA and a 1.21% increase for OA on the RSIVQA dataset). In Fig. 3, we compare the results of RSAdapter and Bi-Modal when using only 10% and 20% of the data on LR and HR datasets. RSAdapter outperforms Bi-Modal on both LR and HR datasets, except for the AA on 10% of LR. This discrepancy is due to the class imbalance in the LR dataset, where the Rural/Urban category has a significant impact on the final AA. In Fig. 4, we also present the results of RSAdapter on the RSIVQA dataset across various training subsets, different categories, AA and OA. In conclusion, our method demonstrates the ability to achieve satisfactory results even with limited data.
@@
IV-D5 Different Bottleneck Dimension Size
TABLE IX: Effect of bottleneck size $d^{\prime}$ of RSAdapters on the LR test set. Results are reported as percentage (%). Best results are highlighted in bold.
@@ IV-D5 Different Bottleneck Dimension Size
The bottleneck dimension size of RSAdapters is a crucial hyperparameter that affects both the number of model parameters and tunable parameters. To investigate its impact on model performance, we select $d^{\prime}$ from the set ${32,64,128,192,256,384}$. Table IX presents the results of the model on the LR test set for different values of $d^{\prime}$. Upon analyzing the results, we observe that the model performs best when $d^{\prime}$ is set to 192. However, as $d^{\prime}$ increases, the model’s parameter count also rises, but there is no corresponding improvement in performance. This suggests that the increase in $d^{\prime}$ may lead to increased optimization complexity, resulting in worse performance. Therefore, selecting the appropriate $d^{\prime}$ is crucial.
IV-D6 Different Position of RSAdapters
In addition to the experiments mentioned above, we also conduct experiments to assess the impact of the insertion positions of RSAdapter, using the same tunable parameters for a fair comparison. Table X provides results for different insertion positions: ’Top’ refers to inserting RSAdapter into the last 6 layers of the transformer, ’Bottom’ refers to inserting RSAdapter into the first 6 layers of the transformer, while ’Even’ and ’Odd’ stand for inserting RSAdapter into the even and odd layers of the transformer, respectively. Remarkably, when RSAdapter is inserted into the lower layers, the results consistently outperform those when inserted into the upper layers. This could be attributed to the fact that there may be slight differences in low-level features between natural images and remote sensing images, leading to significant variations in results when fine-tuning a model pre-trained on natural images for remote sensing images.
TABLE X: Effect of different positions of RSAdapters on the LR test set. Results are reported as percentage (%). Best results are highlighted in bold.